Income taxes may be ignored for this problem. ABC Company Calculated the NPV of a project and found it to be: The projects life was estimated to be: The required rate of
Answer #1

NPV = Present value of cash flows - Initial Investment

$49,800 = $124,000 * PVAF (12% , 9 years) - Initial Investment

Initial Investment = $124,000 * 5.32825 - $49,800

Initial Investment = $660,703 - $49,800

Initial Investment = $610,903

Cross check

Year 1 2 3 4 5 6 7 8 9 Cash flows Present Value Factor Present Value A B = (1/R)^n C = A*B R = 12% -6,10,903.00 1 -6,10,903.0